Two filament lamp A and B takes 0.8Ampere respectively when connected across 110 volt supply.Determine the value of current when they are connected in series

1
Expert's answer
2021-12-07T20:56:52-0500

I=0.8Amp

V=110 volt

Lamp A resistance = R1

Lamp B resistance=R2

equlalent resistance

"\frac{1}{R_{eq}}=\frac{1}{R_1}+\frac{1}{R_2}"

"R_{eq}=\frac{R_1\times R_2}{R_1+R_2}"

Puut value

Now each resistance current flo 0.8amp then both resistance are equal

"R_1=R_2=R"

"R_{eq}=\frac{R\times R}{R+R}=\frac{R}{2}"

Ohm's law

"V=IR_{eq}"

"R_{eq}=\frac{V}{I}"

"R=\frac{2V}{I}"

Put value

"R=\frac{2\times110}{0.8}=275 \Omega"

Now both filament are series combination


"R_{eq}=R_1+R_2"

"R_{eq}=275+275=550\Omega"

Now calculate current

Series combination current floweach bulb is same

"I=\frac{V}{R_{eq}}"

"I=\frac{110}{550}=0.2Amp"

Each bulb current flow

"I_1=I_2=0.2Amp"